a cylindrical diving bell,whose length is 150cm is lowered to the bottom of the tank.the water is found to raise 50cm in the bell. find depth of the tank.atmospheric pressure is 1000cm of water and temperature as constant.

Grade:12

1 Answers

Utkarsh bajpai
6 years ago
 
pressure at surface =1000x150a
pressure in water=100ax(1000-50+h)
P1v1=p2v2
(950+h)100=1000x150
h=1500-950=550